write a slope intercept form of equation of line through (-3, -2), slope= -2/3
Answer by algebrahouse.com(1659)

(-3,-2), m = -2/3
y - y1 = m(x - x1) {point-slope form}
y - (-2) = -2/3[x - (-3)] {substituted into point-slope form}
y + 2 = -2/3(x + 3) {simplified}
y + 2 = (-2/3)x - 2 {used distributive property}
y = (-2/3)x - 4 {subtracted 2 from each side}

For more help from me, visit: www.algebrahouse.com